Patients (20) with high blood pressure were randomly assigned into two groups, a placebo group and a treatment group. The placebo group received conventional treatment while the treatment group received a new drug that is expected to lower blood pressure. After treatment for a couple of months, blood pressures of the patients were measured.


Determine if the new drug had a significant effect on high blood pressure.


-----------------------------------------------------  

    Blood Pressure (systolic)


Placebo Group New Drug Group

-----------------------------------------------------  

90 71

95 79

67 69

120 98

89 91

92 85

100 89

83 75

79 78

85 80


1
Expert's answer
2021-10-18T15:17:35-0400

The hypothesis to be tested for this case is,

"H_0:\mu_1=\mu_2"

"Against"

"H_1:\mu_1\lt\mu_2"

To perform this hypothesis test, we use the "p-value" method. The obtained "p-value" is then compared with the default "\alpha=0.05" since we are not given the level of significance.

Applying the 2 sample t-test in "R", the test can be performed by keying in the following commands,

x=c(90,95,67,120,89,92,100,83,79,85)

y=c(71,79,69,98,91,85,89,75,78,80)

t. test(x, y)

Running these commands gives us the following output,

 Welch Two Sample t-test


data: x and y

t = 1.6081, df = 15.569, p-value = 0.1279

alternative hypothesis: true difference in means is not equal to 0

95 percent confidence interval:

 -2.730473 19.730473

sample estimates:

mean of x mean of y 

   90.0   81.5 

The "p-value" from this output is divided by 2 since the test we are conducting is one sided.

Therefore, "p-value=0.1279\/2=0.06395" and the null hypothesis is rejected if "p-value\lt\alpha" .

Since "p-value=0.06395\gt\alpha=0.05," we fail to reject the null hypothesis and conclude that there is no sufficient evidence to show that the new drug had a significant effect on high blood pressure as compared to placebo at "\alpha=0.05" level of significance.